1. 2H2 + O2 ® 2H2O

a.) How many moles of H2 would be required to produce 8.21 moles of water?

 

          

 

b.) What mass of H2O is formed when H2 reacts with 57 g of O2?

 

1
Expert's answer
2021-03-10T06:58:41-0500

a) "=\frac{8.21 moles H_2O×2 mol H_2}{2 mol H_2O}"


"=8.21mol H_2"


b) The mol ratio between "O_2" and "H_2O" is 1:2, meaning there are 2 moles of water produced for every every 1 mole of Oxygen reacted with Hydrogen.


"=\frac{57g O_2} {32g/mol}=1.78mol O_2"


=2×1.78=3.56mol of water produced.


Since water has a molar Mass of 18g/mol


"=3.56 mol \u00d718g\/mol"


=64.08g "H_2O"